OpenStudy (anonymous):

if xy=28 yz=18 zx=14 and x , y , z>0 the value of x+ y + z is equal to?

OpenStudy (anonymous):

@experimentX @glgan1

OpenStudy (experimentx):

must be some error there xy=28 yx=18

OpenStudy (anonymous):

sorry yz=18

OpenStudy (turingtest):

\[y=\frac {28}x\]\[z=\frac{14}x\]so\[yz=(\frac{28}x)(\frac{14}x)=18\]should work...

OpenStudy (anonymous):

wat u r trying to say plzz help me dude.......................

OpenStudy (turingtest):

I'm sort of thinking out loud...

OpenStudy (turingtest):

\[yz=\frac{392}{x^2}=18\implies x=\frac{14}3\]now you can get the other numbers

OpenStudy (anonymous):

\[ x=\frac{28}{y}\\ z=\frac{18}{y}\\ 14=\frac{28} y \frac{y}{18} y=6 \]

OpenStudy (anonymous):

When a student ask for help and you try to help, they issue you a warning unless you say some jargon between each step. @open_study1 sent me a pm to help.

OpenStudy (anonymous):